OpenStudy (anonymous):

A capacitor of capacitance 5 uF is connected to a 3.0 V battery. How much electrical potential energy is stored? 1.5 x 10-5 J 1.7 x 10-5 J 6.0 x 10-6 J 2.3 x 10-5 J

OpenStudy (radar):

Use \[w _{c}=(1/2)CE ^{2}\]That will get you close enough.

OpenStudy (shamim):

Given C=5microfarad=5*10^-6F E=3 volt Wc=?